OpenStudy (anonymous):

describe the behavior, as c varies, of the level curve f(x,y) = c for f(x,y) = x^2 + y^2 + 1

OpenStudy (jamesj):

Well, when c = 1, what points do you have? What about when c < 1? And c > 1?

OpenStudy (anonymous):

@JamesJ i'm supposed to do this for c = 0 and 2. do i plug 0 in for x and y, then 2? or 0 for x and 2 for y?

OpenStudy (jamesj):

If c = 0, then you have what? x^2 + y^2 + 1 = 0 i.e., x^2 + y^2 = -1 What points (x,y) satisfy that equation?

OpenStudy (anonymous):

so c= z?

OpenStudy (jamesj):

Set aside z for now. The question is what points (x,y) satisfy the equation f(x,y) = c for given values of c. Start then with c = 0. What values of (x,y) satisfy the equation f(x,y) = 0?

OpenStudy (anonymous):

i'm not quite sure how to get 2 square roots to equal -1...

OpenStudy (jamesj):

Indeed! \[ x^2 \geq 0 \ \ \ \ and \ \ \ \ y^2 \geq 0\] for all x and all y. Hence are there ANY points (x,y) such that f(x,y) = 0?

OpenStudy (anonymous):

....nooo?

OpenStudy (jamesj):

No, none. There are no points (x,y) such that f(x,y) = 0. What about when f(x,y) = 1?